# Foveated Retinotopy Improves Classification and Localization in Convolutional Neural Networks

### v4

@@ Abstract
From falcons detecting prey to humans recognizing faces, many species exhibit remarkably rapid visual abilities in both localization and classification. This is enabled by a specialized retinal region called the fovea, which provides high acuity in the center of vision while maintaining lower resolution in the periphery. This spatial organization, which is preserved along the early visual pathway by retinotopic mapping, is fundamental to biological vision, yet remains underexplored in machine learning. In this paper, we investigate how incorporating foveated retinotopy can benefit deep convolutional neural networks (CNNs) in image classification tasks. By implementing a foveated retinotopic transformation as a preprocessing layer for standard off-the-shelf models and retraining them, we maintain comparable classification accuracy while significantly enhancing the network’s robustness to scale and rotation perturbations. Although this architectural change introduces increased sensitivity to fixation point shifts, we demonstrate that this apparent limitation actually becomes advantageous: variations in classification probabilities across gaze positions create a proxy for a saliency map that effectively facilitates object localization. Our results suggest that foveated retinotopic mapping encodes implicit knowledge about visual object geometry, providing an efficient solution to the visual search problem, a fundamental capability critical for many living species. These findings bridge biological vision principles with artificial neural networks, offering new directions for building more robust and efficient computer vision systems.
Author summary
Many animals, including falcons, hummingbirds, sharks and humans, have a remarkable structure in their eyes called the fovea – a small region that provides sharp, detailed vision at the center of our gaze, functioning as a natural high resolution lens. When we want to see something clearly, we move our eyes to center this high-resolution area directly on it. This efficient design has evolved over millions of years, yet modern artificial intelligence systems process images fundamentally differently, examining each pixel with equal weight. In this research, we explore a compelling question: what if we teach computers to see more like animals do? By implementing a fovea-like mechanism in artificial neural networks, we create systems that not only maintain their object recognition capabilities but also gain significant new advantages. Like biological visual systems, these modified networks become more robust when processing images from different distances and angles. Most intriguingly, they also develop a powerful emergent ability: by sampling different viewpoints, they can effectively search for and locate objects of interest – analogous to how animals scan a scene with their eyes. Our findings demonstrate that borrowing designs from nature can make artificial vision systems not just more biologically plausible, but also more versatile and efficient. This biologically-inspired approach offers promising new directions for computer vision that align with millions of years of evolutionary optimization.
@@
For predators like birds of prey (Mitkus et al., 2018) or sharks (Collin, 2018), efficiently detecting prey is critical for survival. More broadly, visual search —the process by which organisms scan their environment to locate and identify objects of interest— is essential across species. This includes predators like falcons and sharks, as well as herbivores such as howler monkeys, fruit bats, hummingbirds, and parrots. Many of these species share a visual system characterized by foveated retinotopy (Polyak, 1941), where the visual field is represented more densely around a central region. This organization is believed to underpin the efficiency of visual search.
@@
While many species, such as prey animals like rabbits, have a uniform retinal topography that assigns roughly equal importance to all parts of the visual field, humans and many mammals possess a foveated retinotopy. This organization features a sharp disparity in resolution, with a high-acuity central area and lower-resolution peripheral regions. Its center defines the point of fixation. Already at the retina, the mapping is mostly radial and the density of photoreceptors decreases exponentially with eccentricity (Anstis, 1974), defining a so called log-polar mapping. This radial organization of the retina, with highest acuity in the center and decreasing acuity in the periphery, is largely maintained across the different stages of visual processing in the brain, as shown in the human cortex by functional magnetic resonance imaging (fMRI) to map visual responses (Tootell et al., 1998; Dougherty et al., 2003). A first study by Tootell et al. (1998) focused on the phenomenon of spatial attention and sought to elucidate the response exhibited when a bar stimulus is presented in a subdivided visual field. As a result, a foveated retinotopic mapping of the visual field is observed in the early visual areas, specifically V1, V2, and V3. Dougherty et al. (2003) then worked out the representation in the same areas when stimuli varying in eccentricity or angular direction were projected in direct association with the log-polar reference. These studies showed how different regions of visual cortex are activated depending on the position of stimuli in the visual field, confirming previously proposed models.
The precise evolutionary advantage of this foveal organization is still under debate. For example, it is thought to facilitate efficient parallel processing of spatial features in both primates (Nauhaus et al., 2016) and humans (Golomb et al., 2008). It may also facilitate connections within and between processing areas that respect the geometry of the sensory epithelium and minimize global wiring length (Lu et al., 2023). However, it has also been suggested that this is merely an artifact of the scaffolding that operates during development (Weinberg, 1997). Another hypothesis is that a fovea implements prior knowledge of visual function with respect to the species’ range of behaviors, as it allows efficient use of resources by prioritizing high-resolution vision in areas where it matters most (i.e., the center of gaze) while sacrificing some resolution in peripheral vision. Eagles or bottlenose dolphins are remarkable in this regard because they have a dual fovea, one for fine lateral vision and one for high acuity straight ahead vision. This suggests that each retinal topography provides an efficient solution to the visual search problem for the given species within its respective ecological niche. More generally, this raises the problem of how the visual system builds an egocentric model of local space from the parcellated retinal input Martins et al. (2024).
@@
A natural question is what computational advantages these retinotopic visual inputs confer on information processing. Numerous hypotheses have been proposed regarding the role of this nonuniform visual field mapping. One primary explanation is that a foveal input facilitates visual exploration: a retina with a fovea allows efficient visual processing if the eye can actively move and focus attention on specific points of interest. More generally, the fovea is associated with a set of oculomotor behaviors aimed at positioning objects of interest at the center of the retina (saccades, smooth pursuit, vestibulo-ocular reflex, …), thus maximizing access to visual information for these objects. In particular, when we visually explore a scene, our eyes perform saccades, rapid eye movements that move the fovea to successive points of interest. In humans, 2 to 3 times per second on average, corresponding to about $150,000$ movements per day and the typical duration of a saccade is very short, about $200$ milliseconds. In between these saccades, fixations allow precise processing of visual information in a short period of time averaging 200 to 300 milliseconds. This alternation between periods of saccades and fixations forms the basis of our scan path and influences how we perceive and interact with our environment. An interesting hypothesis is that the distribution of photoreceptors in foveated retinotopy reflects the probability that an object of interest is located at the center of the retina, given the dynamic parameters of the oculomotor system (Lewis et al., 2003, 2004), thus allowing efficient visual exploration.
@@
To better understand how visual search works in humans, it is necessary to study the underlying mechanisms that control our eye movements and visual attention. It is influenced by several factors, including the saliency of objects (Itti and Koch, 2001). The systematic study of visual search began with the pioneering work of Yarbus (1961) and Noton and Stark (1971), who introduced the concept of a “scan path” as the trajectory of eye movements produced during visual exploration. In the original experiment conducted by Yarbus, participants were presented with a series of different exploration targets applied to realistic paintings or photographs. The results showed that the scan paths that participants followed were not random, but had a degree of structure that was consistent with their individual goals. Noton and Stark (1971) further improved the methods for recording and analyzing these movements by developing an approach that involves non-invasive monocular measurement of eye movements using the diffuse scleral reflection technique, allowing ecological observation of eye movements during recognition tasks (Noton and Stark, 1971). Since then, eye-tracking studies have shown that our eyes follow predictable patterns that maximize the efficiency of visual information acquisition (Rothkopf et al., 2016). Recent studies have shown that this combination of saccades and foveal input, coupled with an effective point-of-interest detection mechanism, significantly enhances visual acuity (Daucé et al., 2020; Daucé and Perrinet, 2020; Dabane et al., 2022) and supports functions such as the integration of local feature analysis into global perceptual representations. This method provides a promising basis for further investigation of eye movements and the function of foveated retinotopy.
@@
These observations have led to the development of novel models of artificial vision using foveated retinotopy. Sandini and Tagliasco (1980) were among the first to develop such models, drawing inspiration from the structures and functions observed in the human eye. In their study, they propose a model that samples the visual scene as a function of eccentricity. It already demonstrates the contribution of this type of transformation to the compression of visual information, emphasizing the visual information in the center and shrinking visual information in the periphery. The most widely used model to represent this topography is the log-polar model, described in detail by Araujo and Dias (1997). It is organized around two polar coordinates: azimuth (angle) and eccentricity (distance from center) on a logarithmic scale. A log-polar sensor would have a resolution (pixel density) that decays exponentially with distance from the center of the image. Recent developments in neuromorphic computing have also shown that a foveated transformation can be implemented in hardware, allowing real-time processing of visual information (Faramarzi et al., 2024).
@@
The log-polar transformation has important consequences for the representation of the geometry of objects in the image. It produces important distortions of the original image, with a strong enlargement of the objects present in the center and a shrinking of the peripheral ones (see figure 1). In addition, the changes produced by movements of the visual sensor, such as translation, rotation, or zoom, have different consequences in a log-polar reference frame. Any movement of the visual sensor that results in a shift of the center of fixation profoundly alters the distribution of pixels, enlarging the central region and blurring the visual information at the periphery (see Figure 1-E&F). Other camera movements have a more limited effect. In particular, any rotation of the sensor around the central fixation point results in a translation in polar space. Similarly, zooming in or out has the effect of a simple translation in the polar dimension (Araujo and Dias, 1997). The log-polar mapping transforms the rotations and zooms into two independent (orthogonal) displacements, since a rotation only changes the azimuth angle $\theta$, while a zoom affects the logarithmic eccentricity $\rho$ on the radial axis. This corresponds to a unique and important property - zooms and rotations in Cartesian space become vertical and horizontal shifts, respectively, in log-polar space (see the green square in Figure 1-A,D&G).
@@
The log-polar retinotopic mapping was successfully applied in computer vision, especially for template matching (Araujo and Dias, 1997; Sarvaiya et al., 2009; Maiello et al., 2020; Palander and Brandt, 2008) or robotics (Javier Traver and Bernardino, 2010; Antonelli et al., 2015). However, implementing visual processing with retinotopic sensors poses significant challenges in terms of information retrieval: Due to the low peripheral resolution, it is difficult to estimate in advance (before an eye movement is made) which part of the image contains the most relevant information. When analyzing a scene, the eye must infer the regions of interest even before they are positioned at the center of the retina. Two broad categories of approaches can be defined to characterize this anticipatory property of the visual system. The first family of methods (called bottom-up) considers only low-level features to characterize the regions that attract the eye (regions that deviate significantly from the average statistics of the scene). This is the approach proposed, for example, by Itti and Koch (2001). Conversely, top-down approaches use prior knowledge about visual content to direct attention to potentially interesting regions that have not yet been explored (Wolfe, 1994). Only a few works exploit this principle of “predictive coding” in the case of retinotopic sensors; see for example (Najemnik and Geisler, 2005; Daucé, 2018).
@@
Recent studies have also explored the potential of foveal retinotopy; Cao et al. (2021) designed a new LPNet architecture that includes an internal layer using the log-polar reference. They showed that rotation invariance during categorization can be improved by adding a layer that encodes the input image in the log-polar reference. da Costa et al. (2024) used a CornetZ to investigate the reorganization of the intermediate layer when training the network with retinal ganglion cell (RGC) sampling, a method that produces Cartesian images that mimic foveal retinotopy. It is shown that the use of such inputs has an impact on CNN receptive field mapping. Finally, also using a sampling method, Lukanov et al. (2021) shows that this non-uniform distribution may be beneficial when considering the attentional mechanism for implementing saccades, as it could improve the categorization accuracy of the networks. It seems that the community is increasingly interested in foveal retinotopy and its benefits for the accuracy and robustness of CNNs. These promising results pave the way to a better understanding of this organization, but they overlook an important point: they all rely on a Cartesian reconstruction of the foveal retinal distribution instead of focusing on the log-polar reference frame.
@@
Deep learning algorithms have made tremendous progress in recent years. For some visual recognition tasks, such as the ImageNet challenge (Russakovsky et al., 2015), convolutional neural networks (CNNs) have made huge strides in computer vision, so much so that they have now surpassed human accuracy in this task (He et al., 2015). CNNs are particularly well suited to image processing tasks because they allow the operations performed by a neural network to be factorized, while limiting the number of parameters dedicated to each processing layer. Convolutional kernels are adaptive and are trained on many examples for image classification. Although CNNs are trained according to a plasticity rule that is not biologically plausible (Bengio et al., 2016), they nevertheless have properties that are close to some exhibited by biological visual systems (Kubilius et al., 2018), in particular (i) massively parallel processing and (ii) the ability to increase the size of the receptive field (the size of a region of the original image) as a function of layer depth. In addition, the “pooling” operator also allows a CNN to be less sensitive to the position (or displacement) of objects in the image: we speak of translation invariance, which can be achieved by alternating convolutional layers and pooling layers with increasingly larger receptive fields. CNNs also show striking similarities to the ability of humans to quickly classify images, such as whether an image contains an animal or not (Jérémie and Perrinet, 2023).
@@
However, one limitation that remains is their vulnerability to adversarial attacks. Studies have shown that these learned models can be fooled by modifications to images that are imperceptible to humans (Huang et al., 2017). These small distortions cause the algorithms to misclassify examples with high confidence (Szegedy et al., 2013). This vulnerability makes deep networks unstable and unsuitable for use in safety-critical domains such as medicine, autonomous vehicles, or other life-or-death situations, and requires proper care to generate more robust responses (Bashivan et al., 2023). Before deep learning can be relied upon for such applications, researchers must find ways to make these models more resilient to adversarial examples and introduce human-level robustness to ensure that mistakes do not have dangerous consequences in the real world. One emerging hypothesis is that insights from biological neuroscience would be critical to achieving this goal (Mineault et al., 2024).
@@
Despite the growing influence of biological and computational neuroscience on computer vision and machine learning, the contribution of foveated retinotopic mapping to visual processing remains scarce. In particular, it should be highly advantageous to exploit the translational invariance within the feature maps of CNNs as an advantage over the geometry of retinotopic mapping. Indeed, it has been shown that the architecture of any deep learning architecture acts as a “deep prior” to constrain the range of achievable representations, so that at one extreme one could use CNNs without learning (Lempitsky et al., 2018). In this context, the log-polar transformation could be seen as a way to constrain the range of transformations that are possible in the input space, thus providing a more efficient way to learn the underlying structure of the visual world.
@@
Several studies have explored transformations to enhance the robustness of CNNs. For example, Remmelzwaal et al. (2019) demonstrated that log-polar transformations positively impact CNN performance and accuracy in digit discrimination tasks. More recently, Kim et al. (2020) proposed an architecture (CyCNN) that leverages polar coordinate mapping and cylindrical convolutional layers to achieve rotational invariance in image classification tasks, particularly on small datasets. Analogous to these approaches, Esteves et al. showed that spherical transformations can confer rotational invariance in 3D object categorization.
@@
Here, we bridge biological and artificial vision by implementing a log-polar transform which we feed to the input of off-the-shelf deep convolutional neural networks (CNNs), creating a foveated mapping that mimics the spatial organization of biological retinas. Our key contributions are:
  • Despite their effectiveness, we demonstrate that conventional CNNs can exhibit significant vulnerability to basic geometric transformations such as rotations and zooms if not considered correctly, highlighting a fundamental limitation in current architectures.
  • We introduce a biologically-inspired foveated architecture by incorporating a log-polar transformation at the network input, embedding rotation and scale invariance into the model’s structure.
  • We validate our approach using transfer learning on standard off-the-shelf architectures and benchmarks, demonstrating that the benefits of foveated vision can be achieved without compromising classification performance when the object is centered.
  • We show how our foveated architecture leverages existing convolutional layers in novel ways to enhance robustness against geometric transformations, providing insights into the relationship between network architecture and invariance properties.
  • We reveal how the log-polar transformation’s sensitivity to fixation point creates an unexpected advantage: by analyzing the probability at the output of the network, we observe variations across multiple viewpoints, such that our system achieves effective object localization without additional training.
  • We establish connections between computational findings and biological vision, offering new directions for both machine learning architectures and our understanding of natural visual processing. By aligning computational models with neurobiological principles, we aim to improve the performance of artificial systems while also contributing more biologically plausible models to the NeuroAI community.

### v5

@@ Abstract
From falcons spotting preys to humans recognizing faces, rapid visual abilities depend on a foveated retinal organization which delivers high‑acuity central vision while preserving low‑resolution periphery. This organization is conserved along early visual pathways but remains underexplored in machine learning. Here we examine how embedding a foveated retinotopic transformation as a preprocessing layer impacts convolutional neural networks (CNNs) for image classification. By applying a log‑polar mapping to off‑the‑shelf models and retraining them, we retain comparable accuracy while improving robustness to scale and rotation. We show that this architecture becomes highly sensitive to fixation‑point shifts, and that this sensitivity yields a proxy for defining saliency maps that effectively facilitates object localization. Our results show that foveated retinotopy encodes prior geometric knowledge, offering a solution to visual‑search and enhancing both classification and localization. These findings connect biological vision principles with artificial networks, pointing to new, robust and efficient directions for computer‑vision systems.
@@ Introduction
For predators like birds of prey Mitkus2018 or sharks Collin2018, efficiently detecting prey is critical for survival. More broadly, visual search —the process by which organisms scan their environment to locate and identify objects of interest— is essential across species. This includes predators like falcons and sharks, as well as herbivores such as howler monkeys, fruit bats, hummingbirds, and parrots. Many of these species share a visual system characterized by foveated retinotopy Polyak1941, where the visual field is represented more densely around a central region. This organization is believed to underpin the efficiency of visual search.
@@
While many species, such as prey animals like rabbits, have a uniform retinal topography that assigns roughly equal importance to all parts of the visual field, humans and many mammals possess a foveated retinotopy. This organization features a sharp disparity in resolution, with a high-acuity central area and lower-resolution peripheral regions. Its center defines the point of fixation. Already at the retina, the mapping is mostly radial and the density of photoreceptors decreases exponentially with eccentricity Anstis1974, defining a so called log-polar mapping. This radial organization of the retina, with highest acuity in the center and decreasing acuity in the periphery, is largely maintained across the different stages of visual processing in the brain, as shown in the human cortex by functional magnetic resonance imaging (fMRI) to map visual responses Tootell1998; Dougherty2003. A first study by Tootell1998 focused on the phenomenon of spatial attention and sought to elucidate the response exhibited when a bar stimulus is presented in a subdivided visual field. As a result, a foveated retinotopic mapping of the visual field is observed in the early visual areas, specifically V1, V2, and V3. Dougherty2003 then worked out the representation in the same areas when stimuli varying in eccentricity or angular direction were projected in direct association with the log-polar reference. These studies showed how different regions of visual cortex are activated depending on the position of stimuli in the visual field, confirming previously proposed models.
The precise evolutionary advantage of this foveal organization is still under debate. For example, it is thought to facilitate efficient parallel processing of spatial features in both primates nauhaus_efficient_2016 and humans golomb_native_2008. It may also facilitate connections within and between processing areas that respect the geometry of the sensory epithelium and minimize global wiring length Lu2023. However, it has also been suggested that this is merely an artifact of the scaffolding that operates during development Weinberg1997. Another hypothesis is that a fovea implements prior knowledge of visual function with respect to the species’ range of behaviors, as it allows efficient use of resources by prioritizing high-resolution vision in areas where it matters most (i.e., the center of gaze) while sacrificing some resolution in peripheral vision. Eagles or bottlenose dolphins are remarkable in this regard because they have a dual fovea, one for fine lateral vision and one for high acuity straight ahead vision. This suggests that each retinal topography provides an efficient solution to the visual search problem for the given species within its respective ecological niche. More generally, this raises the problem of how the visual system builds an egocentric model of local space from the parcellated retinal input Martins2024.
@@
A natural question is what computational advantages these retinotopic visual inputs confer on information processing. Numerous hypotheses have been proposed regarding the role of this nonuniform visual field mapping. One primary explanation is that a foveal input facilitates visual exploration: a retina with a fovea allows efficient visual processing if the eye can actively move and focus attention on specific points of interest. More generally, the fovea is associated with a set of oculomotor behaviors aimed at positioning objects of interest at the center of the retina (saccades, smooth pursuit, vestibulo-ocular reflex, …), thus maximizing access to visual information for these objects. In particular, when we visually explore a scene, our eyes perform saccades, rapid eye movements that move the fovea to successive points of interest. In humans, 2 to 3 times per second on average, corresponding to about $150,000$ movements per day and the typical duration of a saccade is very short, about $200$ milliseconds. In between these saccades, fixations allow precise processing of visual information in a short period of time averaging 200 to 300 milliseconds. This alternation between periods of saccades and fixations forms the basis of our scan path and influences how we perceive and interact with our environment. An interesting hypothesis is that the distribution of photoreceptors in foveated retinotopy reflects the probability that an object of interest is located at the center of the retina, given the dynamic parameters of the oculomotor system lewis2003distribution; lewis2004understanding, thus allowing efficient visual exploration.
@@
To better understand how visual search works in humans, it is necessary to study the underlying mechanisms that control our eye movements and visual attention. It is influenced by several factors, including the saliency of objects itti_koch2005. The systematic study of visual search began with the pioneering work of Yarbus1961 and noton_scanpaths_1971, who introduced the concept of a “scan path” as the trajectory of eye movements produced during visual exploration. In the original experiment conducted by Yarbus, participants were presented with a series of different exploration targets applied to realistic paintings or photographs. The results showed that the scan paths that participants followed were not random, but had a degree of structure that was consistent with their individual goals. noton_scanpaths_1971 further improved the methods for recording and analyzing these movements by developing an approach that involves non-invasive monocular measurement of eye movements using the diffuse scleral reflection technique, allowing ecological observation of eye movements during recognition tasks noton_scanpaths_1971. Since then, eye-tracking studies have shown that our eyes follow predictable patterns that maximize the efficiency of visual information acquisition Rothkopf2016. Recent studies have shown that this combination of saccades and foveal input, coupled with an effective point-of-interest detection mechanism, significantly enhances visual acuity Dauce2020; Dauce2020a; dabane2022you and supports functions such as the integration of local feature analysis into global perceptual representations. This method provides a promising basis for further investigation of eye movements and the function of foveated retinotopy.
@@
These observations have led to the development of novel models of artificial vision using foveated retinotopy. Sandini1980 were among the first to develop such models, drawing inspiration from the structures and functions observed in the human eye. In their study, they propose a model that samples the visual scene as a function of eccentricity. It already demonstrates the contribution of this type of transformation to the compression of visual information, emphasizing the visual information in the center and shrinking visual information in the periphery. The most widely used model to represent this topography is the log-polar model, described in detail by Araujo1997. It is organized around two polar coordinates: azimuth (angle) and eccentricity (distance from center) on a logarithmic scale. A log-polar sensor would have a resolution (pixel density) that decays exponentially with distance from the center of the image. Recent developments in neuromorphic computing have also shown that a foveated transformation can be implemented in hardware, allowing real-time processing of visual information Faramarzi2024.
@@
The log-polar transformation has important consequences for the representation of the geometry of objects in the image. It produces important distortions of the original image, with a strong enlargement of the objects present in the center and a shrinking of the peripheral ones (see figure 1). In addition, the changes produced by movements of the visual sensor, such as translation, rotation, or zoom, have different consequences in a log-polar reference frame. Any movement of the visual sensor that results in a shift of the center of fixation profoundly alters the distribution of pixels, enlarging the central region and blurring the visual information at the periphery (see Figure 1-E&F). Other camera movements have a more limited effect. In particular, any rotation of the sensor around the central fixation point results in a translation in polar space. Similarly, zooming in or out has the effect of a simple translation in the polar dimension Araujo1997. The log-polar mapping transforms the rotations and zooms into two independent (orthogonal) displacements, since a rotation only changes the azimuth angle $\theta$, while a zoom affects the logarithmic eccentricity $\rho$ on the radial axis. This corresponds to a unique and important property - zooms and rotations in Cartesian space become vertical and horizontal shifts, respectively, in log-polar space (see the green square in Figure 1-A,D&G).
@@
The log-polar retinotopic mapping was successfully applied in computer vision, especially for template matching Araujo1997; Sarvaiya2009; Maiello2020; Palander2008 or robotics JavierTraver2010; Antonelli2015. However, implementing visual processing with retinotopic sensors poses significant challenges in terms of information retrieval: Due to the low peripheral resolution, it is difficult to estimate in advance (before an eye movement is made) which part of the image contains the most relevant information. When analyzing a scene, the eye must infer the regions of interest even before they are positioned at the center of the retina. Two broad categories of approaches can be defined to characterize this anticipatory property of the visual system. The first family of methods (called bottom-up) considers only low-level features to characterize the regions that attract the eye (regions that deviate significantly from the average statistics of the scene). This is the approach proposed, for example, by itti_koch2005. Conversely, top-down approaches use prior knowledge about visual content to direct attention to potentially interesting regions that have not yet been explored wolfe1994guided. Only a few works exploit this principle of “predictive coding” in the case of retinotopic sensors; see for example najemnik2005optimal; dauce2018active.
@@
Recent studies have also explored the potential of foveal retinotopy; Cao2021 designed a new LPNet architecture that includes an internal layer using the log-polar reference. They showed that rotation invariance during categorization can be improved by adding a layer that encodes the input image in the log-polar reference. da_costa_convolutional_2024 used a CornetZ to investigate the reorganization of the intermediate layer when training the network with retinal ganglion cell (RGC) sampling, a method that produces Cartesian images that mimic foveal retinotopy. It is shown that the use of such inputs has an impact on CNN receptive field mapping. Finally, also using a sampling method, Lukanov2021 shows that this non-uniform distribution may be beneficial when considering the attentional mechanism for implementing saccades, as it could improve the categorization accuracy of the networks. It seems that the community is increasingly interested in foveal retinotopy and its benefits for the accuracy and robustness of CNNs. These promising results pave the way to a better understanding of this organization, but they overlook an important point: they all rely on a Cartesian reconstruction of the foveal retinal distribution instead of focusing on the log-polar reference frame.
@@
Deep learning algorithms have made tremendous progress in recent years. For some visual recognition tasks, such as the ImageNet challenge Russakovsky2015, convolutional neural networks (CNNs) have made huge strides in computer vision, so much so that they have now surpassed human accuracy in this task He2015wrn. CNNs are particularly well suited to image processing tasks because they allow the operations performed by a neural network to be factorized, while limiting the number of parameters dedicated to each processing layer. Convolutional kernels are adaptive and are trained on many examples for image classification. Although CNNs are trained according to a plasticity rule that is not biologically plausible Bengio2016, they nevertheless have properties that are close to some exhibited by biological visual systems Kubilius2018, in particular (i) massively parallel processing and (ii) the ability to increase the size of the receptive field (the size of a region of the original image) as a function of layer depth. In addition, the “pooling” operator also allows a CNN to be less sensitive to the position (or displacement) of objects in the image: we speak of translation invariance, which can be achieved by alternating convolutional layers and pooling layers with increasingly larger receptive fields. CNNs also show striking similarities to the ability of humans to quickly classify images, such as whether an image contains an animal or not Jeremie2023.
@@
However, one limitation that remains is their vulnerability to adversarial attacks. Studies have shown that these learned models can be fooled by modifications to images that are imperceptible to humans Huang2017. These small distortions cause the algorithms to misclassify examples with high confidence szegedy2013intriguing. This vulnerability makes deep networks unstable and unsuitable for use in safety-critical domains such as medicine, autonomous vehicles, or other life-or-death situations, and requires proper care to generate more robust responses Bashivan2023. Before deep learning can be relied upon for such applications, researchers must find ways to make these models more resilient to adversarial examples and introduce human-level robustness to ensure that mistakes do not have dangerous consequences in the real world. One emerging hypothesis is that insights from biological neuroscience would be critical to achieving this goal Mineault2024.
@@
Despite the growing influence of biological and computational neuroscience on computer vision and machine learning, the contribution of foveated retinotopic mapping to visual processing remains scarce. In particular, it should be highly advantageous to exploit the translational invariance within the feature maps of CNNs as an advantage over the geometry of retinotopic mapping. Indeed, it has been shown that the architecture of any deep learning architecture acts as a “deep prior” to constrain the range of achievable representations, so that at one extreme one could use CNNs without learning Lempitsky2018. In this context, the log-polar transformation could be seen as a way to constrain the range of transformations that are possible in the input space, thus providing a more efficient way to learn the underlying structure of the visual world.
@@
Several studies have explored transformations to enhance the robustness of CNNs. For example, Remmelzwaal2019 demonstrated that log-polar transformations positively impact CNN performance and accuracy in digit discrimination tasks. More recently, Kim2020 proposed an architecture (CyCNN) that leverages polar coordinate mapping and cylindrical convolutional layers to achieve rotational invariance in image classification tasks, particularly on small datasets. Analogous to these approaches, esteves_learning_2018 showed that spherical transformations can confer rotational invariance in 3D object categorization.
@@
Here, we bridge biological and artificial vision by implementing a log-polar transform which we feed to the input of off-the-shelf deep convolutional neural networks (CNNs), creating a foveated mapping that mimics the spatial organization of biological retinas. Our key contributions are the following.
  • First, despite their effectiveness, we demonstrate that conventional CNNs can exhibit significant vulnerability to basic geometric transformations such as rotations and zooms if not considered correctly, highlighting a fundamental limitation in current architectures.
  • Then, we introduce a biologically-inspired foveated architecture by incorporating a log-polar transformation at the network input, embedding rotation and scale invariance into the model’s structure.
  • Next, we validate our approach using transfer learning on standard off-the-shelf architectures and benchmarks, demonstrating that the benefits of foveated vision can be achieved without compromising classification performance when the object is centered.
  • Following that point, we show how our foveated architecture leverages existing convolutional layers in novel ways to enhance robustness against geometric transformations, providing insights into the relationship between network architecture and invariance properties.
  • Moreover, we reveal how the log-polar transformation’s sensitivity to fixation point creates an unexpected advantage: by analyzing the probability at the output of the network, we observe variations across multiple viewpoints, such that our system achieves effective object localization without additional training.
  • Finally, we establish connections between computational findings and biological vision, offering new directions for both machine learning architectures and our understanding of natural visual processing. By aligning computational models with neurobiological principles, we aim to improve the performance of artificial systems while also contributing more biologically plausible models to the NeuroAI community.